# Blending Is All You Need: Cheaper, Better Alternative to Trillion-Parameters LLM

### 09 Jan 2024 v2

@@ Abstract
In conversational AI research, there’s a noticeable trend towards developing models with a larger number of parameters, exemplified by models like ChatGPT. While these expansive models tend to generate increasingly better chat responses, they demand significant computational resources and memory. This study explores a pertinent question: Can a combination of smaller models collaboratively achieve comparable or enhanced performance relative to a singular large model? We introduce an approach termed Blending, a straightforward yet effective method of integrating multiple chat AIs. Our empirical evidence suggests that when specific smaller models are synergistically blended, they can potentially outperform or match the capabilities of much larger counterparts. For instance, integrating just three models of moderate size (6B/13B paramaeters) can rival or even surpass the performance metrics of a substantially larger model like ChatGPT (175B+ paramaters). This hypothesis is rigorously tested using A/B testing methodologies with a large user base on the Chai research platform over a span of thirty days. The findings underscore the potential of the Blended strategy as a viable approach for enhancing chat AI efficacy without a corresponding surge in computational demands.  [^1] All trained models are provided at https://huggingface.co/ChaiML.
@@ 3.3 Blended
3:  $u_{k}\leftarrow\text{user's current input turn}$
@@ 5 Experiments
Blend (13,6,6B)GPT3.5 (175B)Vicuna+ (13B)ChaiLLM (6B)05050100100Improvement Over Control %EngagementRetention
@@ 5.1 Experimental Set Up
00.20.20.40.40.60.60.80.8110.20.20.40.40.60.60.80.8111.21.21.41.4GPT3.5 (175B)Blend (13,6,6B)Vicuna+ (13B)ChaiLLM (6B)Relative Inference Speed (1/FLOPs)Improvement Over BaselineEngagement vs Inference Speed
@@
00.20.20.40.40.60.60.80.81100.20.20.40.40.60.60.80.8GPT3.5 (175B)Blend (13,6,6B)Vicuna+ (13B)ChaiLLM (6B)Relative Inference Speed (1/FLOPs)Improvement Over BaselineRetention vs Inference Speed

### 23 Jan 2024 v3

@@ Abstract
In conversational AI research, there’s a noticeable trend towards developing models with a larger number of parameters, exemplified by models like ChatGPT. While these expansive models tend to generate increasingly better chat responses, they demand significant computational resources and memory. This study explores a pertinent question: Can a combination of smaller models collaboratively achieve comparable or enhanced performance relative to a singular large model? We introduce an approach termed Blending, a straightforward yet effective method of integrating multiple chat AIs. Our empirical evidence suggests that when specific smaller models are synergistically blended, they can potentially outperform or match the capabilities of much larger counterparts. For instance, integrating just three models of moderate size (6B/13B parameters) can rival or even surpass the performance metrics of a substantially larger model like ChatGPT (175B+ parameters). This hypothesis is rigorously tested using A/B testing methodologies with a large user base on the Chai research platform over a span of thirty days. The findings underscore the potential of the Blended strategy as a viable approach for enhancing chat AI efficacy without a corresponding surge in computational demands.  [^1] All trained models are provided at https://huggingface.co/ChaiML.
@@ 3.3 Blended
[…]
@@ 5 Experiments
Blend (13,6,6B)GPT3.5 (175B)Vicuna+ (13B)ChaiLLM (6B)0020202020404040406060606080808080100100100100120120120120Improvement Over Control %EngagementRetention
@@ 5.1 Experimental Set Up
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811110.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811111.21.21.21.21.41.41.41.4GPT3.5 (175B)Blend (13,6,6B)Vicuna+ (13B)ChaiLLM (6B)Relative Inference Speed (1/FLOPs)Improvement Over BaselineEngagement vs Inference Speed
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.8GPT3.5 (175B)Blend (13,6,6B)Vicuna+ (13B)ChaiLLM (6B)Relative Inference Speed (1/FLOPs)Improvement Over BaselineRetention vs Inference Speed
Figure 2: User Engagement